OpenStudy (karlaltr):

Use the appropriate compound angle formula to determine exact values Tan (pi/3-pi/6)

OpenStudy (karlaltr):

Tan(pi/3)-tan(pi/6) -------------------------- 1-tan(pi/3)tan(pi/6)

OpenStudy (holsteremission):

Looks like you've got it... Do you know the values of \(\tan\dfrac{\pi}{3}\) and \(\tan\dfrac\pi6\)?

OpenStudy (karlaltr):

Im not sure how to find those values :/

OpenStudy (mathmale):

Those 2 values of the tangent function are important to know, and thus worth the time and effort to learn. pi/6 = 30 degrees. tan pi/6 = tan 30 deg = [sin 30] / [cos 30] = 1/2 over sqrt(3)/2. What 's your final result here, after reduction of the last fraction?

OpenStudy (holsteremission):

^That should be enough to finish the problem. Not sure why you're asked to use the compound angle formula, since \(\dfrac\pi3-\dfrac\pi6=\dfrac\pi6\), but I suppose you should still practice and know the formula...